Skip to main content

IT Status Alerts

All Closed Alerts

Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245 246 247 248 249 250 251 252 253 254 255 256 257 258 259 260 261 262

5240 results found, displaying results 561 - 581.

Scheduled Service Change: Printing (Net-Print): Apply bug fixes, Sunday, August 6th beginning 6am

Start:
2017-08-06 06:00:00
Status:
Closed
Brief Description:
Apply a bug fix update to Net-Print PaperCut application server and printer servers. This bug fix resolves slow performance of the admin web pages.
more details

Scheduled Service Change: Upgrade the remainder of the Managed FireWall (MFW) service.

Start:
2017-08-06 05:00:00
Status:
Closed
Brief Description:
Upgrade the remainder of the Managed Firewall Service. This includes upgrading FortiManager (FMG) and MFW FortiGates (FGT). The Rhodes and Humphreys firewalls were upgraded successfully on July 16th and this will upgrade the rest.
more details

Performance Issue: Slowdown of email through e-lists

Start:
2017-08-03 14:27:00
Status:
Closed
Brief Description:
A slowdown of email going through e-lists hosted on the main e-list server occurred around 10:15am. Additionally during that time the administrative web was responding slowly. As of 12:10pm everything is responding normally.
Current Status:
As of 12:10pm everything is responding normally.
more details

Unplanned Outage: Network connectivity in Filter Plant

Start:
2017-08-03 01:26:00
Status:
Closed
Brief Description:
The network switch Filter1-2500 is currently down. No network connectivity for Filter Plant Bldg 5210 310 Caldwell Rd.
Current Status:
The Filter1-2500 switch came backup at 11:19AM. Network and WiFi connectivity has been restored to Filter Plant Bldg.
more details

Performance Issue: F5 load balancer

Start:
2017-08-03 07:40:00
Status:
Closed
Brief Description:
An issue was encountered with the F5 load balancer during this mornings scheduled maintenance.
Current Status:
Configs all synced and host are up and passing traffic.
more details

Scheduled Service Change: Move CNE head end to a new appliance.

Start:
2017-08-03 05:00:00
Status:
Closed
Brief Description:
On Thursday, August 3rd between 5am and 7am, CIT Network Engineering will move the Cornell Network Extension Service head end to a new appliance. Customers of CNE will experience a 10 minute disruption in connectivity to Cornell campus networks during the
more details

Scheduled Service Change: PeopleSoft: Bi-weekly migration of online and batch objects. 8/3/2017

Start:
2017-08-03 05:00:00
Status:
Closed
Brief Description:
PeopleSoft: Bi-weekly migration of online and batch objects.
more details

Performance Issue: Confluence Prod turned off anonymous access

Start:
2017-07-30 11:00:00
Status:
Closed
Brief Description:
Confluence upgrade was successful this morning however we are experiencing some performance issues after. We have temporarily turned off anonymous access and pointing those users to the test Confluence site for content.
Current Status:
CIT turned on anonymous access in confluence on Sunday at 3:05PM.
more details

Performance Issue: E-list Delays

Start:
2017-07-31 06:45:00
Status:
Closed
Brief Description:
The main e-list (Lyris) server started having issues at 6:45am July 31, 2017. The web interface was restored about 8:45am. E-list email is still encountering delays.
Current Status:
The problem has been resolved and as of 11:25am all pending e-list email has been delivered. The main e-list server is caught up and back in normal operation.
more details

Scheduled Service Change: SVC (San Volume Controller) maintenance in Rhodes Hall

Start:
2017-07-30 06:00:00
Status:
Closed
Brief Description:
On Sunday 07/30/2017 from 6:00am-noon CIT will be performing IBM SVC (SAN Volume Controller) maintenance in Rhodes Hall.
more details

Scheduled Service Change: KFS, eShop, Cynergy, Travel and Cynergy unavailable--new code deployment

Start:
2017-07-30 06:00:00
Status:
Closed
Brief Description:
KFS, eShop, Cynergy, Cynergy Applications and Travel will be down for new code deployment.
more details

Scheduled Service Change: Video on Demand Download Feature--May Impact Prev. Uploaded Media

Start:
2017-07-30 09:00:00
Status:
Closed
Brief Description:
On Sunday July 30, 2017, we are broadening the options available to let owners allow others to download copies of their media. The change may impact existing users if they had already found and enabled the download feature.
more details

Scheduled Service Change: Confluence AWS (Wikis): Upgrade on Sunday, July 30th, 6 am

Start:
2017-07-30 06:00:00
Status:
Closed
Brief Description:
Confluence AWS (Wikis) application Upgrade in order to bring it back under vendor support and an upgrade to Apache.
more details

Scheduled Service Change: Kronos: planned outage for semi-annual database server OS patching

Start:
2017-07-29 04:00:00
Status:
Closed
Brief Description:
Kronos: planned outage for semi-annual database server OS patching
more details

COECIS Alert: Doing a 2nd COECIS alert test

Start:
2017-07-28 00:00:00
Status:
Closed
Brief Description:
Checking on other features
more details

COECIS Alert: Trying a third COECIS alert

Start:
2017-07-28 00:00:00
Status:
Closed
Brief Description:
One more test
more details

COECIS Alert: Testing CoE CIS alerts

Start:
2017-07-28 00:00:00
Status:
Closed
Brief Description:
Adding an update to test that part
more details

Scheduled Service Change: TeamForge/SVN/SourceForge Prod Linux OS additional update

Start:
2017-07-25 04:00:00
Status:
Closed
Brief Description:
TeamForge/SVN/SourceForge Prod Linux OS additional update
more details

Unplanned Outage: The Network Switches at American Indian House and 10 Sisson Pl are down

Start:
2017-07-24 19:39:00
Status:
Closed
Brief Description:
The Network Switches at American Indian House and 10 Sisson Pl are down due to NYSEG power outage.
Current Status:
NYSEG power was restored after midnight 01:15AM.
more details

COECIS Alert: Testing COECIS alert

Start:
2017-07-25 00:00:00
Status:
Closed
Brief Description:
Testing functionality of COECIS alerts
more details
Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245 246 247 248 249 250 251 252 253 254 255 256 257 258 259 260 261 262